# Multimodal Generative Flows for LHC Jets

### v2

@@ 1 Introduction
The Large Hadron Collider (LHC) at CERN produces billions of proton–proton collisions per second, reconstructed into final-state particles by multi-layered detectors. Among the many objects emerging from hadronic collisions, jets—collimated sprays of localized high-energy particles—play a central role in both QCD studies and searches for new physics. As particle clouds, jets have become a testbed for modern generative models. For example, they can be used to learn background distributions directly from data for resonant anomaly detection, bypassing the limitations of imperfect Monte Carlo simulations and avoiding reliance on high-level, hand-crafted observables [Buhmann et al. (2024)]. These models can also act as tractable high-dimensional density estimators, offering a principled framework for evaluating how closely jet classifiers approach the theoretical optimum [Geuskens et al. (2024)].
Dynamics-based frameworks such as diffusion [Sohl-Dickstein et al. (2015); Song et al. (2020)] and flow-matching [Albergo & Vanden-Eijnden (2022); Lipman et al. (2022)] have recently set new benchmarks for jet generation and now underpin state-of-the-art foundation models for particle physics [Mikuni & Nachman (2025)]. However, these methods operate solely on continuous spaces. This is inadequate for jets, where each particle carries both continuous kinematics and categorical attributes like charge and flavor. The use of de-quantization methods or modeling these modalities separately risks distorting physically meaningful correlations. Accurate modeling therefore requires a framework that jointly treats both continuous and discrete modalities. Multimodal flows have recently gained attention in other scientific fields [Campbell et al. (2024)], see related work in App A.
We propose a multimodal generative framework that integrates continuous flow-matching with a continuous-time Markov jump process for the discrete dynamics. This yields a unified probabilistic path over hybrid spaces capable of generating both kinematics and quantum numbers for jet constituents. Trained on CMS Open Data, our model accurately reproduces the kinematics, substructure and flavor content of real world jets. We show that our model, when equipped with a multimodal particle transformer architecture, with mode-specific and fused encoders can produce state of the art results on the AspenOpenJets dataset introduced by Amram et al. (2024).
@@ 2 CMS Open Data
In this work we are interested in training generative models on real LHC jets. We use the recently released AspenOpenJets (AOJ) dataset derived from 13 TeV proton-proton collisions recorded by CMS in 2016. Each jet is represented as a particle-cloud $\mathbf{z}=\{z^{d}\}_{d=1}^{D}$ with up to $D=150$ constituents, where each particle is described by (continuous) kinematic features in hadronic coordinates and a (categorical) flavor token: $z^{d}\equiv(x^{d},\,k^{d})\in\mathbb{R}^{3}\otimes\mathcal{F}$ with $x^{d}\in(p_{T},\,\Delta\eta,\,\Delta\phi)$ and $k^{d}\in\mathcal{F}=\{\gamma,\,h^{0},\,h^{-},\,h^{+},\,e^{-},\,e^{+},\,\mu^{-},\,\mu^{+}\}$. Here $p_{T}=\sqrt{p_{x}^{2}+p_{y}^{2}}$ is the transverse momentum, and $\Delta\eta$, $\Delta\phi$ are coordinates relative to the jet axis. The possible flavors are: photons ($\gamma$), electrically charged hadrons ($h^{-}\,,\ h^{+}$), electrically neutral hadrons ($h^{0}$), electrons ($e^{-}$), positrons ($e^{+}$), muons $(\mu^{-})$ and anti-muons ($\mu^{+}$), corresponding to particle species and charges reconstructed by the CMS tracking system and calorimeters. The flavor composition is strongly imbalanced: photons and charged hadrons make up $\sim$90% of all constituents (in almost equal proportions), neutral hadrons $\sim$10%, while leptons occur only at the per-mille level.
@@ 3 Multimodal flows for particle clouds
where the posterior follows $\pi_{t}(\bm{z}_{0},\bm{z}_{1}|\bm{z})={p}_{t}(\bm{x}|\bm{x}_{0},\bm{x}_{1}){q}_{t}(\bm{k}|\bm{k}_{0},\bm{k}_{1})\mu(\bm{z}_{0})\nu(\bm{z}_{1})\,/\,P_{t}(\bm{z})$ from Bayes theorem. Although each component evolves independently under the conditional process, non-trivial correlations between particles and their respective modalities emerge for the generative process (2).
@@
The next step is to specify the conditional dynamics over $\mathbb{R}^{3}\otimes\mathcal{F}$. For the continuous modality, the standard choice is to take a uniform flow that transports source data into target data through straight paths with constant velocities [Liu et al. (2022)]. For the discrete modality we propose a multivariate generalization of the random telegraph process [Gardiner (2010)], a continuous-time Markov process originally used to model noise in binary communication channels. As shown in App. B.2, the resulting velocity field and jump rate matrix are given by:
@@
We train using the flow-matching mean-square error (MSE) loss for the particle kinematics and the cross-entropy (CE) loss for the particle flavor tokens. Following the approach of Kendall et al. (2018), these two objectives are combined into a single weighted loss:
@@
In the expectation, time is drawn uniformly $t\sim\mathcal{U}[0,1]$ over the unit interval, pairs of source and target points drawn from the coupling $(\bm{z}_{0},\bm{z}_{1})\sim\mu\otimes\nu$ and $\bm{z}_{t}\sim P_{t}(\cdot|\bm{z}_{0},\bm{z}_{1})$. In contrast to Kendall et al. (2018), where the uncertainty weights $\sigma_{i}$ are fixed trainable scalars, we promote them to time-dependent functions $\sigma^{i}_{t}$. In practice, we parametrize the weights via $\sigma_{t}^{i}=\exp(-w_{t}^{i})$, where $w_{t}$ is the output of an uncertainty network that we discard during inference. This allows the relative weighting between modalities to adapt dynamically along the generative path, enabling the model to prioritize different objectives at different stages of the evolution. Besides improving the convergence of the training, this formulation eliminates the need for manually tuning loss weights.
@@
To learn the continuous and discrete modalities together, we approximate the conditional velocity field and the posterior classifier function with a single permutation equivariant neural network $u^{\theta}_{t}\otimes h^{\theta}_{t}$. The overall architecture, depicted in Fig. 1, has three components: 1) two mode-specific encoders, 2) a fused encoder, and 3) two task-specific heads. All three encoders consist of non-causal particle transformers [Qu et al. (2022)] with stacked multi-head self-attention blocks. The regressor head predicts the continuous-valued vector field $u^{\theta}_{t}$ for the MSE loss, while the classifier head outputs the logits $h^{\theta}_{t}$ for the CE loss. Fore more details on the architecture see App. C.
@@
Once the model is trained, new samples can be generated from a source input $\bm{z}_{0}$ by simulating the associated dynamics to the probability flow equation (1). This entails numerically solving the joint dynamics for both the continuous and discrete components. For the continuous dynamics we solve the ordinary differential equation $\dot{\bm{x}}_{t}=\bm{u}_{t}^{\theta}(\bm{x}_{t})$ with Euler’s method using discrete time-steps of size $\Delta t$. The discrete dynamics of the flavor tokens follows the multivariate telegraph process governed by the rate $\mathbf{W}^{\theta}_{t}$ derived from the trained posterior $\pi_{t}^{\theta}={\tt softmax}(h^{\theta}_{t}/T)$. Here we have introduced a temperature scaling hyperparameter $T$ [Guo et al. (2017)] that helps improving sampling quality. To efficiently simulate this stochastic process, we employ $\tau$-leaping [Gillespie (2001); Campbell et al. (2022)], a well-established approximation method widely used in chemical reaction kinetics. More details on the sampling algorithm can be found in App. D.
@@ 4 Experiments
In this section we demonstrate that our transformer-based multimodal flow (MMF) is capable of generating the particle kinematics and flavor composition of real-world jets from CMS open data. Since our work consists of a proof-of-concept, we do not attempt a full optimization of the architectures and only train our models on a subset of the AspenOpenJets dataset. Our training dataset consists of $1.25$ million AOJ jets, with $1$ million jets for training and 250250K jets for validation. For the source data $\bm{z}_{0}=\{z_{0}^{d}\}_{d=1}^{D}$, we draw point-cloud noise from the distribution $z^{d}_{0}=(x^{d}_{0},\,k^{d}_{0})\sim\nu\equiv\mathcal{N}(0,\mathbf{1})\,\otimes\,{\mathcal{U}}(p)$ where $\mathcal{N}$ denotes the standard Gaussian over $\mathbb{R}^{3}$ and $\mathcal{U}$ represents the uniform categorical distribution over the token space $\mathcal{F}$ with probability parameter $p=1/S$. We compare our proposed model to EPiC-FM, a permutation equivariant flow-matching model with deep-sets architecture designed for particle cloud data that achieves state-of-the-art results on simulated jets [Buhmann et al. (2023b, a); Birk et al. (2023)]. Our training setup is described in App. E. We generate samples with 270K jets using a temperature scaling of $T=0.85$ and solve the dynamics with $\Delta t=0.001$ time-step size. To quantify the performance of the continuous component of our generator we compute the jet-level kinematics $(p_{T},m,\eta,\phi)$ and the $N$-subjettiness ratios $\tau_{21}$, $\tau_{32}$ [Thaler & Van Tilburg (2011)] as probes for jet substructure. For the discrete component we examine the particle flavor multiplicities $N^{k}$, $k\in{\mathcal{F}}$. Finally, to test the model capacity in reproducing cross-modal correlations we compute the jet charge $\mathcal{Q}=\sum_{i\in\rm jet}Q_{i}{p_{T}}_{i}/p_{T}^{\rm jet}$, a non-trivial aggregate of the particles electric charge $Q_{i}$ and transverse momentum ${p_{T}}_{i}$. These distributions, and the corresponding Wasserstein-1 distances to the AOJ truth, are provided in Fig. 2 and Table 1, respectively.
@@ Appendix A Related Work
Generative models that independently handle discrete and continuous variables have been explored since the inception of research in both denoising and diffusion frameworks, with various methodologies attempting to relate these approaches. Winkler et al. (2024) show that a limit for Ehrenfest processes defined on discrete spaces converges to an Ornstein-Uhlenbeck dynamic. Further extending the theory, Ren et al. (2024) establish a stochastic integral formulation of discrete diffusion models, generalizing the Poisson random measure through a Lévy-type integral. The prevalence of multimodal flows in the natural sciences has spurred additional multimodal solutions akin to our work [Lin et al. (2025)]. Purely discrete flows were developed by Gat et al. (2024), while in protein design, Campbell et al. (2024) introduced a fully multimodal methodology. Their approach differs from ours in selecting conditional paths in an ad hoc manner, directly interpolating within probability space. Conversely, we introduce a more general strategy allowing diverse probabilistic paths by leveraging Markov bridges, as presented by Chopin et al. (2023). Finally, a general extension of flow matching theory to a broader family of stochastic process generators was proposed by Holderrieth et al. (2024); Lipman et al. (2024). Our method can be viewed as a special instance of this broader framework, where the probability path is derived via a Markov bridge applied to a specific type of reference process.
@@ Appendix B Generative modeling with bridge processes
In our formulation, we closely follow the flow matching methodology of Albergo & Vanden-Eijnden (2022); Lipman et al. (2022) and start by constructing probability paths that follow a prescribed continuity equation. We then construct conditional probability paths with the help of reference process with known generators that enables the construction of bridges between point samples, ensuring transformations from $\bm{z}_{0}$ to $\bm{z}_{1}$. In the context of particle clouds for jets, we will assume that $\bm{z}=(\bm{x},\bm{k})$ is a collection of continuous ($\bm{x}$) and discrete ($\bm{k}$) vectorial random variables. And in the following we will show the construction for $\bm{x}$ and $\bm{k}$ separately.
@@ B.1 Continuous random variables
Now one can prove [Albergo & Vanden-Eijnden (2022); Lipman et al. (2022)] that the desired target vector field, can be obtained from the conditional by:
@@ B.2 Discrete random variables
We are now poised with the task of constructing a conditional probability path with tractable rates, that achieve conditions (21). Since we are dealing with discrete variables, direct interpolation in data space as achieved by equation (14) is impossible. We will follow instead Fitzsimmons et al. (1992) and construct the conditional probability path as a Markov Bridge.
We assume access to a reference process characterized by the rate $\bm{R}_{t}(\bm{k},\bm{j})$. For this process, we also assume we have closed-form solutions to its corresponding master equation Eq. (19), expressed in terms of the conditional probability $\tilde{q}_{s|t}(\bm{k}_{1}|\bm{k})$ of being in state $\bm{k}_{1}$ at time $s$ when the state was $\bm{k}$ at time $t$. Following Fitzsimmons et al. (1992), we can now construct a Markov bridge that satisfies (21) with a tractable conditional rate given by:
@@
where the expectation runs over $t\sim\mathcal{U}(0,1)$, $(\bm{k}_{0},\bm{k}_{1})\sim\mu\times\nu$ and $\bm{k}_{t}\sim q_{t}(\cdot|\bm{k}_{0},\bm{k}_{1})$. Similar to the square loss problem (18), one can show that that the minimizer of the cross entropy loss is given by the proper conditional (posterior) probability as one can relate the cross entropy loss to the regression problem [Bishop & Nasrabadi (2006)] (it suffices to see the regression problem over the one hot encoding expression of the likelihood). In the following we impose further conditions on our reference process such as to make our computations more tractable.

### v3

@@ 1 Introduction
The Large Hadron Collider (LHC) at CERN produces billions of proton–proton collisions per second, reconstructed into final-state particles by multi-layered detectors. Among the many objects emerging from hadronic collisions, jets—collimated sprays of localized high-energy particles—play a central role in both QCD studies and searches for new physics. As particle clouds, jets have become a testbed for modern generative models. For example, they can be used to learn background distributions directly from data for resonant anomaly detection, bypassing the limitations of imperfect Monte Carlo simulations and avoiding reliance on high-level, hand-crafted observables [Buhmann:2023acn]. These models can also act as tractable high-dimensional density estimators, offering a principled framework for evaluating how closely jet classifiers approach the theoretical optimum [geuskens2024fundamental].
Dynamics-based frameworks such as diffusion [sohl2015deep; song2020score] and flow-matching [albergo2022building; lipman2022flow] have recently set new benchmarks for jet generation and now underpin state-of-the-art foundation models for particle physics [Mikuni:2024qsr]. However, these methods operate solely on continuous spaces. This is inadequate for jets, where each particle carries both continuous kinematics and categorical attributes like charge and flavor. The use of de-quantization methods or modeling these modalities separately risks distorting physically meaningful correlations. Accurate modeling therefore requires a framework that jointly treats both continuous and discrete modalities. Multimodal flows have recently gained attention in other scientific fields [campbell2024generative], see related work in App A.
We propose a multimodal generative framework that integrates continuous flow-matching with a continuous-time Markov jump process for the discrete dynamics. This yields a unified probabilistic path over hybrid spaces capable of generating both kinematics and quantum numbers for jet constituents. Trained on CMS Open Data, our model accurately reproduces the kinematics, substructure and flavor content of real world jets. We show that our model, when equipped with a multimodal particle transformer architecture, with mode-specific and fused encoders can produce state of the art results on the AspenOpenJets dataset introduced by Amram:2024fjg.
@@ 2 CMS Open Data
Figure 1: Visualizations of 10 CMS jets in the $(\eta,\phi)$-plane taken from the AspenOpenJets dataset. Each particle cloud is centered around the jet axis. The size of each marker is proportional to the particle’s transverse momentum ($p_{T}$), while the shape and color encode electric charge and flavor, respectively. The upper right corner of each panel indicates the total number of constituents in the respective jet.
In this work we are interested in training generative models on real LHC jets. We use the recently released AspenOpenJets (AOJ) dataset derived from 13 TeV proton-proton collisions recorded by CMS in 2016. Each jet is represented as a particle-cloud $\mathbf{z}=\{z^{d}\}_{d=1}^{D}$ with up to $D=150$ constituents, where each particle is described by (continuous) kinematic features in hadronic coordinates and a (categorical) flavor token: $z^{d}\equiv(x^{d},\,k^{d})\in\mathbb{R}^{3}\otimes\mathcal{F}$ with $x^{d}\in(p_{T},\,\Delta\eta,\,\Delta\phi)$ and $k^{d}\in\mathcal{F}=\{\gamma,\,h^{0},\,h^{-},\,h^{+},\,e^{-},\,e^{+},\,\mu^{-},\,\mu^{+}\}$. Here $p_{T}=\sqrt{p_{x}^{2}+p_{y}^{2}}$ is the transverse momentum, and $\Delta\eta$, $\Delta\phi$ are coordinates relative to the jet axis. The possible flavors are: photons ($\gamma$), electrically charged hadrons ($h^{-}\,,\ h^{+}$), electrically neutral hadrons ($h^{0}$), electrons ($e^{-}$), positrons ($e^{+}$), muons $(\mu^{-})$ and anti-muons ($\mu^{+}$), corresponding to particle species and charges reconstructed by the CMS tracking system and calorimeters. Representative particle clouds from the AOJ dataset are shown in Fig. 1. The flavor composition is strongly imbalanced: photons and charged hadrons make up $\sim$90% of all constituents (in almost equal proportions), neutral hadrons $\sim$10%, while leptons occur only at the per-mille level.
@@ 3 Multimodal flows for particle clouds
where the posterior follows
$\pi_{t}(\bm{z}_{0},\bm{z}_{1}|\bm{z})={p}_{t}(\bm{x}|\bm{x}_{0},\bm{x}_{1}){q}_{t}(\bm{k}|\bm{k}_{0},\bm{k}_{1})\mu(\bm{z}_{0})\nu(\bm{z}_{1})\,/\,P_{t}(\bm{z})$ (5)
from Bayes theorem. Although each component evolves independently under the conditional process, non-trivial correlations between particles and their respective modalities emerge for the generative process (2).
@@
The next step is to specify the conditional dynamics over $\mathbb{R}^{3}\otimes\mathcal{F}$. For the continuous modality, the standard choice is to take a uniform flow that transports source data into target data through straight paths with constant velocities [liu2022flowstraightfastlearning]. For the discrete modality we propose a multi-state generalization of the random telegraph process [gardiner2010stochastic], a continuous-time Markov process originally used to model noise in binary communication channels (bit-flips). As shown in App. B.2, the resulting velocity field and jump rate matrix are given by:
@@
We train using the flow-matching mean-square error (MSE) loss for the particle kinematics and the cross-entropy (CE) loss for the particle flavor tokens. Following the approach of kendall2018multi, these two objectives are combined into a single weighted loss:
@@
In the expectation, time is drawn uniformly $t\sim\mathcal{U}[0,1]$ over the unit interval, pairs of source and target points drawn from the coupling $(\bm{z}_{0},\bm{z}_{1})\sim\mu\otimes\nu$ and $\bm{z}_{t}\sim P_{t}(\cdot|\bm{z}_{0},\bm{z}_{1})$. In contrast to kendall2018multi, where the uncertainty weights $\sigma_{i}$ are fixed trainable scalars, we promote them to time-dependent functions $\sigma^{i}_{t}$. In practice, we parametrize the weights via $\sigma_{t}^{i}=\exp(-w_{t}^{i})$, where $w_{t}$ is the output of an uncertainty network that we discard during inference. This allows the relative weighting between modalities to adapt dynamically along the generative path, enabling the model to prioritize different objectives at different stages of the evolution. Besides improving the convergence of the training, this formulation eliminates the need for manually tuning loss weights.
@@
To learn the continuous and discrete modalities together, we approximate the conditional velocity field and the posterior classifier function with a single permutation equivariant neural network $u^{\theta}_{t}\otimes h^{\theta}_{t}$. The overall architecture, depicted in Fig. 2, has three components: 1) two mode-specific encoders, 2) a fused encoder, and 3) two task-specific heads. All three encoders consist of non-causal particle transformers [Qu:2022mxj] with stacked multi-head self-attention blocks. The regressor head predicts the continuous-valued vector field $u^{\theta}_{t}$ for the MSE loss, while the classifier head outputs the logits $h^{\theta}_{t}$ for the CE loss. Fore more details on the architecture see App. C.
@@
Once the model is trained, new samples can be generated from a source input $\bm{z}_{0}$ by simulating the associated dynamics to the probability flow equation (1). This entails numerically solving the joint dynamics for both the continuous and discrete components. For the continuous dynamics we solve the ordinary differential equation $\dot{\bm{x}}_{t}=\bm{u}_{t}^{\theta}(\bm{x}_{t})$ with Euler’s method using discrete time-steps of size $\Delta t$. The discrete dynamics of the flavor tokens follows the multivariate telegraph process governed by the rate $\mathbf{W}^{\theta}_{t}$ derived from the trained posterior $\pi_{t}^{\theta}={\tt softmax}(h^{\theta}_{t}/T)$. Here we have introduced a temperature scaling hyperparameter $T$ [guo2017calibration] that helps improving sampling quality. To efficiently simulate this stochastic process, we employ $\tau$-leaping [gillespie2001approximate; NEURIPS2022_b5b52876], a well-established approximation method widely used in chemical reaction kinetics. More details on the sampling algorithm can be found in App. D.
@@ 4 Experiments
Aspen Open Jets
We demonstrate that our transformer-based multimodal flow (MMF) is capable of generating the particle kinematics and flavor composition of real-world jets from CMS open data. Since our work consists of a proof-of-concept, we do not attempt a full optimization of the architectures and only train our models on a subset of the AspenOpenJets dataset. Our training dataset consists of $1.25$ million AOJ jets, with $1$ million jets for training and 250250K jets for validation. For the source data $\bm{z}_{0}=\{z_{0}^{d}\}_{d=1}^{D}$, we draw point-cloud noise from the distribution $z^{d}_{0}=(x^{d}_{0},\,k^{d}_{0})\sim\nu\equiv\mathcal{N}(0,\mathbf{1})\,\otimes\,{\mathcal{U}}(p)$ where $\mathcal{N}$ denotes the standard Gaussian over $\mathbb{R}^{3}$ and $\mathcal{U}$ represents the uniform categorical distribution over the token space $\mathcal{F}$ with probability parameter $p=1/S$. We compare our proposed model to EPiC-FM, a permutation equivariant flow-matching model with deep-sets architecture designed for particle cloud data that achieves state-of-the-art results on simulated jets [Buhmann:2023pmh; Buhmann:2023zgc; Birk:2023efj]. Our training setup is described in App. E. We generate samples with 270K jets using a temperature scaling of $T=0.85$ and solve the dynamics with $\Delta t=0.001$ time-step size. To quantify the performance of the continuous component of our generator we compute the jet-level kinematics $(p_{T},m,\eta,\phi)$ and the $N$-subjettiness ratios $\tau_{21}$, $\tau_{32}$ [Thaler:2010tr] as probes for jet substructure. For the discrete component we examine the particle flavor multiplicities $N^{k}$, $k\in{\mathcal{F}}$. Finally, to test the model capacity in reproducing cross-modal correlations we compute the jet charge
$\mathcal{Q}=\frac{1}{p_{T}^{\rm jet}}\sum_{i\in\rm jet}Q_{i}\,{{p_{T}}_{i}},$ (10)
a non-trivial aggregate of the particles electric charge $Q_{i}$ and transverse momentum ${p_{T}}_{i}$. These distributions, and the corresponding Wasserstein-1 distances to the AOJ truth, are provided in Fig. 3 and Table 1, respectively.
@@ Appendix A Related Work
Generative models that independently handle discrete and continuous variables have been explored since the inception of research in both denoising and diffusion frameworks, with various methodologies attempting to relate these approaches. winkler2024bridging show that a limit for Ehrenfest processes defined on discrete spaces converges to an Ornstein-Uhlenbeck dynamic. Further extending the theory, ren2024discrete establish a stochastic integral formulation of discrete diffusion models, generalizing the Poisson random measure through a Lévy-type integral. The prevalence of multimodal flows in the natural sciences has spurred additional multimodal solutions akin to our work [lin2025tfg]. Purely discrete flows were developed by gat2024discrete, while in protein design, campbell2024generative introduced a fully multimodal methodology. Their approach differs from ours in selecting conditional paths in an ad hoc manner, directly interpolating within probability space. Conversely, we introduce a more general strategy allowing diverse probabilistic paths by leveraging Markov bridges, as presented by chopin2023computational. Finally, a general extension of flow matching theory to a broader family of stochastic process generators was proposed by holderrieth2024generator; lipman2024flow. Our method can be viewed as a special instance of this broader framework, where the probability path is derived via a Markov bridge applied to a specific type of reference process.
@@ Appendix B Generative modeling with bridge processes
In our formulation, we closely follow the flow matching methodology of albergo2022building; lipman2022flow and start by constructing probability paths that follow a prescribed continuity equation. We then construct conditional probability paths with the help of reference process with known generators that enables the construction of bridges between point samples, ensuring transformations from $\bm{z}_{0}$ to $\bm{z}_{1}$. In the context of particle clouds for jets, we will assume that $\bm{z}=(\bm{x},\bm{k})$ is a collection of continuous ($\bm{x}$) and discrete ($\bm{k}$) vectorial random variables. And in the following we will show the construction for $\bm{x}$ and $\bm{k}$ separately.
@@ B.1 Continuous random variables
Now one can prove [albergo2022building; lipman2022flow] that the desired target vector field, can be obtained from the conditional by:
@@ B.2 Discrete random variables
We are now poised with the task of constructing a conditional probability path with tractable rates, that achieve conditions (23). Since we are dealing with discrete variables, direct interpolation in data space as achieved by equation (16) is impossible. We will follow instead fitzsimmons1992markovian and construct the conditional probability path as a Markov Bridge.
We assume access to a reference process characterized by the rate $\bm{R}_{t}(\bm{k},\bm{j})$. For this process, we also assume we have closed-form solutions to its corresponding master equation Eq. (21), expressed in terms of the conditional probability $\tilde{q}_{s|t}(\bm{k}_{1}|\bm{k})$ of being in state $\bm{k}_{1}$ at time $s$ when the state was $\bm{k}$ at time $t$. Following fitzsimmons1992markovian, we can now construct a Markov bridge that satisfies (23) with a tractable conditional rate given by:
@@
where the expectation runs over $t\sim\mathcal{U}(0,1)$, $(\bm{k}_{0},\bm{k}_{1})\sim\mu\times\nu$ and $\bm{k}_{t}\sim q_{t}(\cdot|\bm{k}_{0},\bm{k}_{1})$. Similar to the square loss problem (20), one can show that that the minimizer of the cross entropy loss is given by the proper conditional (posterior) probability as one can relate the cross entropy loss to the regression problem [bishop2006pattern] (it suffices to see the regression problem over the one hot encoding expression of the likelihood). In the following we impose further conditions on our reference process such as to make our computations more tractable.